OpenStudy (anonymous):

help! use the quadratic formula to solve x²+3=-4x

OpenStudy (anonymous):

x^2+4x+3 = 0 (x+1) (x+ 3) = 0 then x = -1 or x = -3

OpenStudy (anonymous):

no no use quadratic formula to solve @gerryliyana

OpenStudy (anonymous):

do you know what the quadratic formula is? or are you asking for that.

OpenStudy (anonymous):

i know what it is but i still dont know how 2 get the answer...

OpenStudy (anonymous):

set up equation so that you have: ax^2+bx+c=0 and use a, b and c correctly in your formula

OpenStudy (anonymous):

x²+4x+3=0?

OpenStudy (anonymous):

so a=?,b=?, and c=?

OpenStudy (anonymous):

yess ur all right

OpenStudy (anonymous):

smart boy!!

OpenStudy (anonymous):

a=o b=4 c=3

OpenStudy (anonymous):

nooo

OpenStudy (anonymous):

a = 1, b=4, c=3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry i wasnt thinking

OpenStudy (anonymous):

so now what?

OpenStudy (anonymous):

use \[x = \frac{ -b \pm \sqrt{(b^{2} - 4ac)} }{ 2a }\]

OpenStudy (anonymous):

substitute all parameter, including a, b, and c

OpenStudy (anonymous):

to solve x²+4x + 3 =0 you should find the value of x, one way to do this by using formula above ...,

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b^{2}-4ac} }{ 2a }\] \[x = \frac{ -4 \pm \sqrt{4^{2}-(4)(1)(3)} }{ 2(1) }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thanks both of you!